OpenStudy (anonymous):

summatiion 1=n+1-m m stand for what?its m=1?

OpenStudy (anonymous):

Requires a different set notation. What you wrote is vague. Summation has 3 different things involved. Something written on top which shows the ending, something at the bottom where it says like i=1 which shows the starting value and an actual function inside the summation like Ai

OpenStudy (anonymous):

\[\sum_{k=m}^{n} 1=n+1-m\]

OpenStudy (anonymous):

this my question actly,n what is m stand for?

OpenStudy (anonymous):

I believe it just means the Summation of 1 at value M to N is equal to n+1-m. It's just saying From the values of M to N it is equal to n+1-m

OpenStudy (anonymous):

can I substitute with 1?

OpenStudy (anonymous):

Nope. That summation is just stating, If you add up all the values of 1 from M to N, it equals n+1-m.

OpenStudy (anonymous):

It's just a general statement that can easily be proven by plugging in a number. So let's say M = 5 and N = 7. 7+1-5 = 3. Summation of 1 (5,7) which is 1+1+1 = 3. (At M=5 it is 1, then all the way up to 7, so that's 5,6,7, so there are 3 "1"

OpenStudy (anonymous):

my equation is this \[12/n \sum_{k=m}^{n}1\]so,=12/n(n+1-m) right then how to solve

OpenStudy (anonymous):

(12/n) (n+1-m). Then just simplify ---> 12(n+1-m)/n ---> (12n+12-12m)/n

OpenStudy (anonymous):

then how to solve by lim\[n approach \infty\]

OpenStudy (anonymous):

For limits in this case since I shall assume M is a constant, if the degree of the unknown "N" is the same (N^1 for denominator and numerator) then you can take the coefficient ratios to find the limit. So 12n/n = 12. The limit as n approaches infinity is 12

OpenStudy (anonymous):

if m is constant then?if limit infinity,m/n right?

OpenStudy (anonymous):

hyunalee....

hero (hero):

Good luck with this

OpenStudy (anonymous):

I don't understand your question. If M is constant then the limit as it approaches infinity is equal to 12. However, if it's not a constant it becomes a little more complicated and I am unsure on how to continue on with this problem.

OpenStudy (anonymous):